OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. -8-5i/6-3i

OpenStudy (anonymous):

-33-54i/45

OpenStudy (alfie):

\[\frac{8-5i}{6-3i} * \frac{6+3i}{6+3i}\] Recall the properties of complex conjugates, and i^2 = -1 and solve ;)

OpenStudy (anonymous):

Alfie thanks see i would have missed that I keep confusing myself on these types of problems then get frustrated and want to throw my computer across the room.

OpenStudy (alfie):

You're welcome, actually they are not too hard, each time you have a division you have to get rid of the i at the denominator so you multiply and divide for the complex conjugate. Then you have to do the products at the numerator, you'll have to recall that i^2 = -1. Now you are good to go. :)

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

when you are done verify with my answer above

OpenStudy (anonymous):

-48-24i-30i-15i^2/36+18i-18i-9i^2 =-48-54i+15/36+9 =33-54i/45